OpenStudy (anonymous):

Multiply. -2x(x^2+6x-9) .... (sorry in advance don't know how to make the 2 show as an exponent)

OpenStudy (johnweldon1993):

just distribute that -2x into the parenthesis

OpenStudy (anonymous):

x^m ( x^n )= x^m+n x^2 (x^2) = x^4

OpenStudy (johnweldon1993):

just incase you aren't certain of what that means...you multiply each term in the parenthesis by the -2x i.e -2x times x² -2x times 6x -2x times -9

OpenStudy (anonymous):

so how many x are you multiplying? x^2 = x * x x^ 3 = x* x* x x^ 4 = x*x*x*x